# Toward Learning POMDPs Beyond Full-Rank Actions and State Observability

### v1

@@ A.2 Proof of Proposition 1
What remains is to show that we can apply $P$ to recover the POMDP initial belief, observation matrices, transition matrices, and final summing vector from the linear PSR models. Let $a\in\mathcal{A}$ and $o\in\mathcal{O}$. Following Eq. 7, we know that $M^{ao}=A_{\mathit{hists}^{-ao,:}}^{\dagger}\cdot\mathcal{H}_{\mathit{hists}^{ao},:}\cdot{V^{T}}^{\dagger}$. If we apply $P$ as a similarity transform, we find that
@@ A.3 Proof of Lemma 1
Consider the event that $w\in\mathbb{S}^{|\mathcal{A}_{\mathit{full}}|\cdot|\mathcal{O}|}$, such that $\Lambda_{ii}=\sum_{k}^{|\mathcal{A}_{\mathit{full}}|\cdot|\mathcal{O}|}w_{k}O^{(ao)_{k}}_{i}i$ and $\Lambda_{jj}=\sum_{k}^{|\mathcal{A}_{\mathit{full}}|\cdot|\mathcal{O}|}w_{k}O^{(ao)_{k}}_{j}j$ are equivalent. Written in terms of the notation introduced above, we have that $\left\langle a-b,w\right\rangle=0$. This means that $w$ must be contained in the hyperplane $H=\{x\in\mathbb{R}^{|\mathcal{A}_{\mathit{full}}|\cdot|\mathcal{O}|}:\left\langle x,a-b\right\rangle=0\}$, which also passes through the origin. We recognize that $H\cap\mathbb{S}^{|\mathcal{A}_{\mathit{full}}|\cdot|\mathcal{O}|}$ is a $|\mathcal{A}_{\mathit{full}}|\cdot|\mathcal{O}|-2$-dimensional submanifold (a lower-dimensional sphere) of $\mathcal{S}^{|\mathcal{A}_{\mathit{full}}|\cdot|\mathcal{O}|}$. We know that the measure of this submanifold under the induced uniform measure on $\mathbb{S}^{|\mathcal{A}_{\mathit{full}}|\cdot|\mathcal{O}|-1}$ from the Lebesgue measure on $\mathbb{R}^{|\mathcal{A}_{\mathit{full}}|\cdot|\mathcal{O}|-1}$ is zero (Lee2012). Thus, the probability of sampling $w$ so that $\Lambda_{ii}=\Lambda_{jj}$ is zero as well. Therefore, the complement of this event, that $\Lambda_{ii}\neq\Lambda_{jj}$, must occur with probability one. $\square$
Remark.
@@ A.4 Proof that Similarity Transform is Recovered up to Block-Diagonal Matrix
where the blocks $Q_{i}\in\mathbb{R}^{|S_{i}|\times|S_{i}|}$ are nonsingular w.p. 1, where $|S_{i}|$ is the $i^{\text{th}}$ partition in the permuted index ordering.
@@
A.5 Proof of Theorem 1
Section 4.3 claims that applying the matrix $P^{\prime}R\operatorname{diag}(R^{T}P^{\prime-1}m_{\infty})$ is a similarity transformation $\tilde{P}$ that satisfies the implication of Theorem 1. As a reminder, $P^{\prime}$ are the eigenvectors from the eigendecomposition of the matrix in Eq. 18 and $R$ is a random block-diagonal rotation matrix with the same block structure as $PP^{\prime-1}=Q$ (Lemma 4), whose blocks are distributed over the Haar measure over the corresponding copy of $SO(n)$.
@@ A.5 Proof of Theorem 1
If we unpack the block structure of $Q$, $R$, and $\operatorname{diag}(R^{T}Q^{-1}m_{\infty})$ in Equations 21 and 23, we find
@@
where $R_{i}$ and $Q_{i}$ are the blocks associated with the full-rank observability partition $S_{i}$.
First, we must justify that the relations expressed in Equations 21-25 are well-defined. We already know $R$ and $Q$ are nonsingular. We must then show all entries of the vector $R^{T}Q^{-1}\mathbf{1}$ are nonzero to allow for the existence of $\operatorname{diag}(R^{T}Q^{-1}\mathbf{1})$. This fact is a consequence of known properties of the Haar measure over special orthogonal matrices (meckesRandomMatrix2019, Section 1.2). Fix a full-rank observability partition $S_{i}$. It is known that corresponding rotation matrix blocks $R_{i}^{T}$ and $R_{i}$ are identically distributed with respect to the Haar measure on $SO(|S_{i}|)$ (meckesRandomMatrix2019, pg. 18). Furthermore, since $Q_{i}$ is nonsingular, we know that $Q_{i}^{-1}\mathbf{1}$ is not the zero vector. Thus, it is also known that the random vector $R_{i}^{T}\cdot(Q_{i}^{-1}[\mathbf{1}]_{S_{i}})$ is uniformly distributed over the $(|S_{i}|-1)$-sphere with radius $\left\lVert Q_{i}^{-1}[\mathbf{1}]_{S_{i}}\right\rVert_{2}$ (meckesRandomMatrix2019, pg. 19-20, 26). By the same argument discussed in the proof of Lemma 1, the entries of $R_{i}^{T}Q_{i}^{-1}[\mathbf{1}]_{S_{i}}$ must be nonzero with probability one. By taking a union bound over all full-rank observability partitions, all entries of $R^{T}Q^{-1}\mathbf{1}$ must be nonzero with probability one as well.
What remains is to prove the correctness of the relations 13-15 in Theorem 1. The expression that we obtain $\tilde{P}^{-1}m_{\infty}=\mathbf{1}$ is immediate from Equations 23 and 25. Next, justify Equation 13. Fix an full-rank observability partition $S_{i}$. Then
@@
$=[b_{\pi}^{T}]_{S_{i}}Q_{i}R_{i}\operatorname{diag}(R_{i}^{T}Q_{i}^{-1}[\mathbf{1}]_{S_{i}})\cdot\operatorname{diag}(R_{i}^{T}Q_{i}^{-1}\mathbf{[}1]_{S_{i}})^{-1}R_{i}^{T}Q_{i}^{-1}\cdot[\mathbf{1}]_{S_{i}}$
@@
The second line applies Equations 24 and 25. The proof for Equation 14 is nearly the same, and can be reached by deriving an analagous expression to Equation 24 by first multiplying out the corresponding sequence of matrices $\tilde{P}^{-1}M^{ao}P$ and unpacking the block structure of $Q$ and $R$ again. $\square$
@@ A.6 Proof of Full-Rank Transition Claim
Proof.
We first observe that the proof is immediate if $p=0$, so we focus on the case for $p\in(0,1)$. Suppose, for the sake of contradition, that there exists $v\in\mathbb{R}^{n}$ such that matrix-vector product $\left(pT+(1-p)I\right)v=0$. This must be true if and only if
@@
We prove this claim by induction on the number of rows and columns. As the base-case, we take a $1\times 1$ “matrix” $\begin{pmatrix}1\end{pmatrix}$. The eigenvalue of this matrix is unity. Next, we assume that the claim holds for $m\times m$ matrices with rows of all zeros except for a single one. Suppose we have a $(m+1)\times(m+1)$ matrix $T^{\prime}$ of the same structure. If $T^{\prime}$ is a permutation matrix, then we know its eigenvalues are roots of unity (artinAlgebra2011), so suppose that $T^{\prime}$ is not a permutation matrix. Then $T^{\prime}$ must have at least one columns that is all zeros. We then examine the characteristic polynomial $\phi(T^{\prime})$. Without loss of generality, assume that column is the first column of the matrix. Then, we can write out the expression for the characteristic polynomial $\phi(T^{\prime})$:
@@
$\phi(T^{\prime})=\det\left(T^{\prime}-\lambda I\right)=\det\left(\begin{array}[]{c | c}-\lambda&\cdots\\ \hline\cr 0&T^{\prime}_{2:,2:}-\lambda I\\ \end{array}\right)=-\lambda\cdot\det(T^{\prime}_{2:,2:}-\lambda I)=-\lambda\cdot\phi(T^{\prime}_{2:,2:})$
@@
where $T^{\prime}_{2:,2:}$ is the $m\times m$ submatrix of $T^{\prime}$ that omits the first row and column of $T^{\prime}$. This submatrix is also a matrix with all zeros for every row except for a single one, since we eliminated a column of only zeros from $T^{\prime}$. Thus, we know that the eigenvalues of $T^{\prime}$ are 0 and the eigenvalues of $T^{\prime}_{2:,2:}$, which, by the induction hypothesis, are also zero and roots of unity. ∎
Appendix B Additional Algorithmic Details
@@ B.1 Parameters Introduced for Finite Data
While the transition and observation likelihoods computed from the data will converge to the values true values asymptotically, approximation error prevents us from directly reading the parameter estimates as probabilities (guoPACRL2016; azizzadenesheliReinforcementLearning2016). Before using the learned model we project all parameters back to probability distributions via quadratic programming by minimizing the $L_{2}$ norm.
Algorithm pseudocode can be found in Alg. 1. We note that for all experiments, while it is theoretically correct to construct a block-diagonal rotation matrix $R$ by the full-rank observability partition as stated above, we find in practice it is sufficient to multiply by fully-dense random rotation matrix $R^{\prime}$ after computing the initial SVD (line 5). This modification uses $AR^{\prime}$ and $R^{\prime T}V$ to compute the linear PSR and takes $R=I$ instead at line 20. We still require a $\tau_{obs}$ parameter to compute the partition-level transition errors in Figures 3 and 4.

### v2

@@ A.2 Proof of Proposition 1
What remains is to show that we can apply $P$ to recover the POMDP initial belief, diagonal observation matrices, transition matrices, and final summing vector from the linear PSR models. Let $a\in\mathcal{A}$ and $o\in\mathcal{O}$. Following Eq. 5, we know that $M^{ao}=A_{\mathit{hists}^{-ao,:}}^{\dagger}\cdot\mathcal{H}_{\mathit{hists}^{ao},:}\cdot{V^{T}}^{\dagger}$. If we apply $P$ as a similarity transform, we find that
@@ A.3 Proof of Lemma 1
Consider the event that $w\in\mathbb{S}^{|\mathcal{A}_{\mathit{full}}|\cdot|\mathcal{O}|}$, such that $\Lambda_{ii}=\sum_{k}^{|\mathcal{A}_{\mathit{full}}|\cdot|\mathcal{O}|}w_{k}O^{(ao)_{k}}_{i}i$ and $\Lambda_{jj}=\sum_{k}^{|\mathcal{A}_{\mathit{full}}|\cdot|\mathcal{O}|}w_{k}O^{(ao)_{k}}_{j}j$ are equivalent. Written in terms of the notation introduced above, we have that $\left\langle a-b,w\right\rangle=0$. This means that $w$ must be contained in the hyperplane $H=\{x\in\operatorname{\mathbb{R}}^{|\mathcal{A}_{\mathit{full}}|\cdot|\mathcal{O}|}:\left\langle x,a-b\right\rangle=0\}$, which also passes through the origin. We recognize that $H\cap\mathbb{S}^{|\mathcal{A}_{\mathit{full}}|\cdot|\mathcal{O}|}$ is a $|\mathcal{A}_{\mathit{full}}|\cdot|\mathcal{O}|-2$-dimensional submanifold (a lower-dimensional sphere) of $\mathcal{S}^{|\mathcal{A}_{\mathit{full}}|\cdot|\mathcal{O}|}$. We know that the measure of this submanifold under the induced uniform measure on $\mathbb{S}^{|\mathcal{A}_{\mathit{full}}|\cdot|\mathcal{O}|-1}$ from the Lebesgue measure on $\mathbb{R}^{|\mathcal{A}_{\mathit{full}}|\cdot|\mathcal{O}|-1}$ is zero [Lee2012]. Thus, the probability of sampling $w$ so that $\Lambda_{ii}=\Lambda_{jj}$ is zero as well. Therefore, the complement of this event, that $\Lambda_{ii}\neq\Lambda_{jj}$, must occur with probability one. $\square$
Remark 1
@@ A.4 Proof that Similarity Transform is Recovered up to Block-Diagonal Matrix
where the blocks $Q_{i}\in\operatorname{\mathbb{R}}^{|S_{i}|\times|S_{i}|}$ are nonsingular w.p. 1, where $|S_{i}|$ is the $i^{\text{th}}$ partition in the permuted index ordering.
@@
A.5 Proof of Theorem 4.1
Section 4.3 claims that applying the matrix $P^{\prime}R\operatorname{diag}(R^{T}P^{\prime-1}m_{\infty})$ is a similarity transformation $\tilde{P}$ that satisfies the implication of Theorem 4.1. As a reminder, $P^{\prime}$ are the eigenvectors from the eigendecomposition of the matrix in Eq. 16 and $R$ is a random block-diagonal rotation matrix with the same block structure as $PP^{\prime-1}=Q$ (Lemma 4), whose blocks are distributed over the Haar measure over the corresponding copy of $SO(n)$.
@@
If we unpack the block structure of $Q$, $R$, and $\operatorname{diag}(R^{T}Q^{-1}m_{\infty})$ in Equations 20 and 22, we find
@@
where $R_{i}$ and $Q_{i}$ are the blocks associated with the full-rank observability partition $S_{i}$, and $U]_{\mathcal{I}}$ represents the values of the vector or matrix-valued quantity $U$ indexed by a set of indices $\mathcal{I}$.
First, we must justify that the relations expressed in Equations 20-24 are well-defined. We already know $R$ and $Q$ are nonsingular. We must then show all entries of the vector $R^{T}Q^{-1}\mathbf{1}$ are nonzero to allow for the existence of $\operatorname{diag}(R^{T}Q^{-1}\mathbf{1})$. This fact is a consequence of known properties of the Haar measure over special orthogonal matrices [meckesRandomMatrix2019, Section 1.2]. Fix a full-rank observability partition $S_{i}$. It is known that corresponding rotation matrix blocks $R_{i}^{T}$ and $R_{i}$ are identically distributed with respect to the Haar measure on $SO(|S_{i}|)$ [meckesRandomMatrix2019, pg. 18]. Furthermore, since $Q_{i}$ is nonsingular, we know that $Q_{i}^{-1}\mathbf{1}$ is not the zero vector. Thus, it is also known that the random vector $R_{i}^{T}\cdot(Q_{i}^{-1}[\mathbf{1}]_{S_{i}})$ is uniformly distributed over the $(|S_{i}|-1)$-sphere with radius $\left\lVert Q_{i}^{-1}[\mathbf{1}]_{S_{i}}\right\rVert_{2}$ [meckesRandomMatrix2019, pg. 19-20, 26]. By the same argument discussed in the proof of Lemma 1, the entries of $R_{i}^{T}Q_{i}^{-1}[\mathbf{1}]_{S_{i}}$ must be nonzero with probability one. By taking a union bound over all full-rank observability partitions, all entries of $R^{T}Q^{-1}\mathbf{1}$ must be nonzero with probability one as well.
What remains is to prove the correctness of the relations 11-13 in Theorem 4.1. The expression that we obtain $\tilde{P}^{-1}m_{\infty}=\mathbf{1}$ is immediate from Equations 22 and 24. First, we justify Equation 11. Fix an full-rank observability partition $S_{i}$. Then
@@
$=[b_{\pi}^{T}]_{S_{i}}Q_{i}R_{i}\operatorname{diag}(R_{i}^{T}Q_{i}^{-1}[\mathbf{1}]_{S_{i}})\cdot\operatorname{diag}(R_{i}^{T}Q_{i}^{-1}[\mathbf{1}]_{S_{i}})^{-1}R_{i}^{T}Q_{i}^{-1}\cdot[\mathbf{1}]_{S_{i}}$
@@
The second line applies Equations 23 and 24. The proof for Equation 12 is nearly the same, and can be reached by deriving an analagous expression to Equation 23 by first multiplying out the corresponding sequence of matrices $\tilde{P}^{-1}M^{ao}P$ and unpacking the block structure of $Q$ and $R$ again. $\square$
@@ A.6 Proof of Full-Rank Transition Claim
Proof
We first observe that the proof is immediate if $p=0$, so we focus on the case for $p\in(0,1)$. Suppose, for the sake of contradition, that there exists $v\in\operatorname{\mathbb{R}}^{n}$ such that matrix-vector product $\left(pT+(1-p)I\right)v=0$. This must be true if and only if
@@
We prove this claim by induction on the number of rows and columns. As the base-case, we take a $1\times 1$ “matrix" $\begin{pmatrix}1\end{pmatrix}$. The eigenvalue of this matrix is unity. Next, we assume that the claim holds for $m\times m$ matrices with rows of all zeros except for a single one. Suppose we have a $(m+1)\times(m+1)$ matrix $T^{\prime}$ of the same structure. If $T^{\prime}$ is a permutation matrix, then we know its eigenvalues are roots of unity [dingWhenMatrix2014], so suppose that $T^{\prime}$ is not a permutation matrix. Then $T^{\prime}$ must have at least one columns that is all zeros. We then examine the characteristic polynomial $\phi(T^{\prime})$. Without loss of generality, assume that column is the first column of the matrix. Then, we can write out the expression for the characteristic polynomial $\phi(T^{\prime})$:
$\phi(T^{\prime})$ $=\det\left(T^{\prime}-\lambda I\right)$
@@
$=\det\left(\begin{array}[]{c | c}-\lambda&\cdots\\ \hline\cr 0&T^{\prime}_{2:,2:}-\lambda I\\ \end{array}\right)$
$=-\lambda\cdot\det(T^{\prime}_{2:,2:}-\lambda I)$
$=-\lambda\cdot\phi(T^{\prime}_{2:,2:})$
@@
where $T^{\prime}_{2:,2:}$ is the $m\times m$ submatrix of $T^{\prime}$ that omits the first row and column of $T^{\prime}$. This submatrix is also a matrix with all zeros for every row except for a single one, since we eliminated a column of only zeros from $T^{\prime}$. Thus, we know that the eigenvalues of $T^{\prime}$ are 0 and the eigenvalues of $T^{\prime}_{2:,2:}$, which, by the induction hypothesis, are also zero and roots of unity. $\square$
A.7 Proof of Theorem 6.1
We first begin by characterizing the stochastic process $\{s_{t}\}_{t=0}^{\infty}$, the marginals of states variables during exploration ($a_{t}\sim\operatorname{Unif}(\mathcal{A})$ i.i.d.). We begin by proving the following lemma:
Lemma 5
The stochastic process $\{s_{t}\}_{t\geq 0}$ is a Markov chain, and its associated transition matrix is
$T=\frac{1}{|\mathcal{A}|}\sum_{a\in\mathcal{A}}T^{a}.$
Proof
We first begin with conditional joint likelihood that we would like to factorize and unmarginalize present and past actions. Because actions are sampled i.i.d., we have that:
$\operatorname{\mathrm{P}}(s_{t},s_{t+2}|s_{t+1})$
$=\sum_{a_{t},a_{t+1}\in\mathcal{A}}\operatorname{\mathrm{P}}(s_{t},s_{t+2}|a_{t},s_{t+1},a_{t+1})\operatorname{\mathrm{P}}(a_{t}|s_{t+1})\operatorname{\mathrm{P}}(a_{t+1}|s_{t+1})$
$=\sum_{a_{t},a_{t+1}\in\mathcal{A}}\operatorname{\mathrm{P}}(s_{t},s_{t+2}|a_{t},s_{t+1},a_{t+1})\operatorname{\mathrm{P}}(a_{t}|s_{t+1})\operatorname{\mathrm{P}}(a_{t+1}|s_{t+1})$
$=\sum_{a_{t},a_{t+1}\in\mathcal{A}}\operatorname{\mathrm{P}}(s_{t}|a_{t},s_{t+1})\operatorname{\mathrm{P}}(s_{t+2}|s_{t+1},a_{t+1})\operatorname{\mathrm{P}}(a_{t}|s_{t+1})\operatorname{\mathrm{P}}(a_{t+1}|s_{t+1})$
$=\sum_{a_{t},a_{t+1}\in\mathcal{A}}\operatorname{\mathrm{P}}(s_{t},a_{t}|s_{t+1})\operatorname{\mathrm{P}}(s_{t+2},a_{t+1}|s_{t+1})$
$=\operatorname{\mathrm{P}}(s_{t}|s_{t+1})\operatorname{\mathrm{P}}(s_{t+2}|s_{t+1})$
The third line applies the fact that $s_{t+2}\perp s_{t}|s_{t+1},a_{t},a_{t+1}$ due to the conditional independencies introduced by the POMDP generative model, and that $a_{t}$ and $a_{t+1}$ are independent of all other variables.
We then observe that the transition likelihood between two states is the averaged transition likelihood across all actions.
$\operatorname{\mathrm{P}}(s_{t+1}|s_{t})=\sum_{a_{t}\in\mathcal{A}}\operatorname{\mathrm{P}}(s_{t+1}|s_{t},a_{t})\operatorname{\mathrm{P}}(a_{t}|s_{t})=\sum_{a_{t}\in\mathcal{A}}\frac{\operatorname{\mathrm{P}}(s_{t+1}|s_{t})}{|\mathcal{A}|}$ (25)
The last equality invokes that $a_{t}\sim\operatorname{Unif}$ i.i.d. Thus, we can conclude from expression 25 that the transition matrix of this Markov chain $T$ can be written as the average of the transition matrices of the POMDP, e.g. $T=\sum_{a\in\mathcal{A}}T^{a}/|\mathcal{A}|$. $\square$
We also observe that the stationary distribution of this Markov chain is $b_{\pi}$, as defined in the proof of Lemma 2, Appendix A.1 and is unique (otherwise, we would violate ergodicity of the Markov chain $\{s_{t},a_{t},o_{t}\}$).
Now, we are ready to prove the main claim of Theorem 6.1. Let the transition and diagonal observation matrices of the original POMDP be $\mathcal{T}=\{T^{a}:a\in\mathcal{A}\}$ and $\mathcal{Z}=\{O^{ao}:a\in\mathcal{A},o\in\mathcal{O}\}$, and suppose that $\mathcal{T}^{\prime}:\{{T^{a}}^{\prime}:a\in\mathcal{A}\}$ and $\mathcal{Z}^{\prime}:\{{O^{ao}}^{\prime}:a\in\mathcal{A},o\in\mathcal{O}\}$ are the transition and diagonal matrices of the POMDP after transformation $P$ is applied. Since $b_{\pi}$ is the unique stationary distribution of Markov chain $\{s_{t}\}_{t\geq 0}$ evolving according to POMDP with transition matrices $\mathcal{T}$, then $b_{\pi}$ must be a unique right-eigenvector of matrix $T=\sum_{a\in\mathcal{A}}T^{a}/|\mathcal{A}|$ with eigenvalue 1 [norrisDiscretetimeMarkov1997]. Therefore, the row-stochastic vector $b_{\pi}^{\prime}=b_{\pi}P^{-1}$ must be a right vector of the transition matrix $T^{\prime}=\sum_{a\in mcA}T^{\prime a}/|\mathcal{A}|$ because
$b_{\pi}^{\prime T}\left(\frac{1}{|\mathcal{A}|}\sum_{a\in\mathcal{A}}T^{\prime a}\right)$
$=b_{\pi}P^{-1}\left(\frac{1}{|\mathcal{A}|}\sum_{a\in\mathcal{A}}PT^{a}P^{-1}\right)$
$=b_{\pi}\left(\frac{1}{|\mathcal{A}|}\sum_{a\in\mathcal{A}}T^{a}\right)P^{-1}$
$=b_{\pi}P^{-1}$
$=b_{\pi}$
From the relation above, we conclude that $b_{\pi}^{\prime}$ must be a unique stationary distribution for the stochastic process $\{s^{\prime}_{t}\}_{t\geq 0}$ that evolves according to POMDP transition matrices $\mathcal{T}^{\prime}$.
Lastly, fix $\mathit{hist}=(a_{1},o_{1},\dots,a_{t},o_{t})$ and $\mathit{test}=(a_{t},o_{t},\dots,a_{n},o_{n})$. We evaluate equations 1 - 3 to relate Hankel matrices $\mathcal{H}$ and $\mathcal{H}^{\prime}$:
$\mathcal{H}_{\mathit{hist},\mathit{test}}$ $=b_{\pi}T^{a_{1}}O^{a_{1}o_{1}}T^{a_{2}}O^{a_{2}o_{2}}\cdots T^{a_{n}}O^{a_{n}o_{n}}\mathbf{1}$
$=b_{\pi}P^{-1}PT^{a_{1}}P^{-1}PO^{a_{1}o_{1}}P^{-1}PT^{a_{2}}P^{-1}\cdots PT^{a_{n}}P^{-1}PO^{a_{n}O_{n}}P^{-1}P\mathbf{1}$
$=b_{\pi}^{\prime}{T^{a_{1}}}^{\prime}{O^{a_{1}o_{1}}}^{\prime}{T^{a_{2}}}^{\prime}{O^{a_{2}o_{2}}}^{\prime}\cdots{T^{a_{n}}}^{\prime}{O^{a_{n}o_{n}}}^{\prime}\mathbf{1}$
$=\mathcal{H}^{\prime}_{\mathit{hist},\mathit{test}}.$
The third line involes the fact that the provided similarity transform in the perturbed sense-float-reset counterexample (Eq. 17) has rows that sum to one. Thus, we conclude that $\mathcal{H}=\mathcal{H}^{\prime}$. $\square$
B Appendix: Additional Algorithmic Details
@@ B.1 Parameters Introduced for Finite Data
While the transition and observation likelihoods computed from the data will converge to the values true values asymptotically, approximation error prevents us from directly reading the parameter estimates as probabilities [guoPACRL2016, azizzadenesheliReinforcementLearning2016]. Before using the learned model we project all parameters back to probability distributions via quadratic programming by minimizing the $L_{2}$ norm.
Algorithm pseudocode can be found in Alg. 1. We note that for all experiments, while it is theoretically correct to construct a block-diagonal rotation matrix $R$ by the full-rank observability partition as stated above, we find in practice it is sufficient to multiply by fully-dense random rotation matrix $R^{\prime}$ after computing the initial SVD (line 5). This modification uses $AR^{\prime}$ and $R^{\prime T}V$ to compute the linear PSR and takes $R=I$ instead at line 20. We still require a $\tau_{obs}$ parameter to compute the partition-level transition errors in Figures 4 and 5.